If t = tan("\theta" /2), show that sin("\theta" )=2t/(1-t2) and cos ("\theta" )=(1-t2)/(1+t2). hence solve the equation cos("\theta" )-2sin("\theta" )=2


1
Expert's answer
2021-10-12T10:37:28-0400

Let "t = \tan(\frac{\theta}2)." Then

"\sin(\theta )=2\sin\frac{\theta}2\cos\frac{\theta}2=2\tan\frac{\theta}2\cos^2\frac{\theta}2
=\frac{2\tan\frac{\theta}2}{\frac{1}{\cos^2\frac{\theta}2}}
=\frac{2\tan\frac{\theta}2}{1+\tan^2\frac{\theta}2}=\frac{2t}{1+t^2}," and

"\cos (\theta )=\cos^2\frac{\theta}2-\sin^2\frac{\theta}2
=\frac{\cos^2\frac{\theta}2-\sin^2\frac{\theta}2}{\cos^2\frac{\theta}2\frac{1}{\cos^2\frac{\theta}2}}
=\frac{1-\tan^2\frac{\theta}2}{1+\tan^2\frac{\theta}2}
=\frac{1-t^2}{1+t^2}."


Let us solve the equation "\cos(\theta )-2\sin(\theta )=2." Let "t = \tan(\frac{\theta}2)." Then we get the equation

"\frac{1-t^2}{1+t^2}-2\frac{2t}{t^2+1}=2," and hence "\frac{1-t^2-4t}{1+t^2}=2." It follows that "1-t^2-4t=2+2t^2," and thus "3t^2+4t+1=0." The last equation is equivalent to"(3t+1)(t+1)=0," and hence has the roots "t_1=-1,\ t_2=-\frac{1}3." Then "\tan(\frac{\theta}2)=-1" or "\tan(\frac{\theta}2)=-\frac{1}3." It follows that "\frac{\theta}2=-\frac{\pi}4+\pi n" or "\frac{\theta}2=-\arctan\frac{1}3+\pi n,\ n\in\mathbb N."

We conclude that the solutions of the equation "\cos(\theta )-2\sin(\theta )=2" are of the form:

"\theta=-\frac{\pi}2+2\pi n" or "\theta=-2\arctan\frac{1}3+2\pi n," where "n\in\mathbb N."